# Qwen2.5-Coder Technical Report

### v1

@@ 6.2 Code Completion
DS-Coder-33B-Base 33B 80.1 89.0 86.8 85.3
Table 7: Performance of different approaches on the HumanEval Infilling Tasks
@@ 6.3 Code Reasoning
CodeQwen1.5-7B-Base 7B 44.8 40.1
Qwen2.5-Coder-7B-Base 7B 56.5 56.0
@@ 6.4 Math Reasoning
Mathematics and coding have always been closely intertwined. Mathematics forms the foundational discipline for coding, while coding serves as a vital tool in mathematical fields. As such, we expect an open and powerful code model to exhibit strong mathematical capabilities as well. To assess Qwen2.5-Coder’s mathematical performance, we selected five popular benchmarks, including MATH (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), MMLU-STEM (Hendrycks et al., 2020) and TheoremQA (Chen et al., 2023). As shown in Table 9, Table 3 highlights Qwen2.5-Coder’s strengths in mathematics, which likely stem from two key factors: first, the model’s strong foundation built on Qwen2.5, and second, the careful mixing of code and mathematical data during training, which has ensured a well-balanced performance across these domains.
@@
CodeQwen1.5-7B-Base 7B 10.6 37.7 39.6 15.8
Qwen2.5-Coder-7B-Base 7B 46.6 83.9 67.6 34.0
@@
Table 9: Performance of various models on four math benchmark, named MATH, GSM8K, MMLU STEM and TheoremQA respectively.
@@
Base Redux
@@
StarCoder2-3B 3B 36.6 37.0
DS-Coder-1.3B-Base 1.3B 25.8 24.5
Qwen2.5-Coder-1.5B 1.5B 53.6 50.9
@@
StarCoder2-7B 7B 38.8 38.6
StarCoder2-15B 15B 64.1 48.8
DS-Coder-6.7B-Base 6.7B 36.4 36.5
DS-Coder-V2-Lite-Base 2.4/16B 60.5 58.3
CodeQwen1.5-7B-Base 7B 40.5 41.2
Qwen2.5-Coder-7B-Base 7B 68.0 66.6
@@
DS-Coder-33B-Base 33B 39.4 38.7
@@
DS-Coder-6.7B-Base 6.7B 36.4 40.2 57.6 53.8
DS-Coder-V2-Lite-Base 2.4/16B 57.3 38.8 72.9 -
CodeQwen1.5-7B-Base 7B 35.7 42.2 59.8 56.0
Qwen2.5-Coder-7B-Base 7B 60.9 50.6 72.9 76.8
@@
Table 11: General performance of different models on four popular general benchmark, ARC-Challenge, TruthfulQA, WinoGrande and HellaSwag.
@@ 6.6 Long-Context Evaluation
Long context capability is crucial for code LLMs, serving as the core skill for understanding repository-level code and becoming a code agent. However, most of current code models still have very limited support for length, which hinders their potential for practical application. Qwen2.5-Coder aims to further advance the progress of open-source code models in long context modeling. To achieve this, we have collected and constructed long sequence code data at the repository level for pre-training. Through careful data proportioning and organization, we have enabled it to support input lengths of up to 128K tokens.
@@
Figure 5: The long context ability of Qwen2.5-Coder, evaluated by Needle in Code.
@@ 7 Evaluation on Instruct Models
Qwen2.5-Coder-1.5B-Instruct https://hf.co/qwen/Qwen/Qwen2.5-Coder-1.5B-Instruct
Qwen2.5-Coder-7B-Instruct https://hf.co/qwen/Qwen/Qwen2.5-Coder-7B-Instruct
CodeQwen1.5-7B-Chat https://huggingface.co/Qwen/CodeQwen1.5-7B-Chat
CodeStral-22B https://hf.co/mistralai/Codestral-22B-v0.1
@@ 7.1 Code Generation
We also evaluated the code generation capabilities of the Qwen2.5-Coder series instruct models using the EvalPlus (Liu et al., 2023) dataset. As illustrated by the experimental results in Table 13, our Qwen2.5-Coder-7B-Instruct model demonstrated superior accuracy, significantly outperforming other models with a comparable number of parameters. Remarkably, it even exceeded the performance of larger models with over 20 billion parameters, such as CodeStral-22B and DS-Coder-33B-Instruct. Notably, Qwen2.5-Coder-7B-Instruct was the only model in our evaluation to surpass an 80% accuracy rate on HumanEval+, achieving an impressive 84.1%.
@@
DS-Coder-1.3B-Instruct 1.3B 65.2 61.6 61.6 52.6 22.8 3.4 9.3
Qwen2.5-Coder-1.5B-Instruct 1.5B 70.7 66.5 69.2 59.4 32.5 6.8 15.7
@@
DS-Coder-6.7B-Instruct 6.7B 78.6 70.7 75.1 66.1 35.5 10.1 20.5
DS-Coder-V2-Lite-Instruct 2.4/16B 81.1 75.0 82.3 68.8 36.8 16.2 24.3
CodeQwen1.5-7B-Chat 7B 86.0 79.3 83.3 71.4 39.6 17.2 20.1
Qwen2.5-Coder-7B-Instruct 7B 88.4 84.1 83.5 71.7 41.0 18.2 37.6
@@
CodeStral-22B 22B 78.1 74.4 73.3 68.2 47.1 20.6 32.9
DS-Coder-33B-Instruct 33B 79.3 68.9 81.2 70.1 46.5 17.6 27.7
@@
DS-Coder-V2-Instruct 21/236B 90.2 82.3 84.8 82.3 83 84.5 79.5 52.5 79.9
Table 14: The performance of different approaches on instruct format MultiPL-E.
@@
The instruct split provided by BigCodeBench (Zhuo et al., 2024) is intended for assessing the code generation abilities of instruct models. We assessed the Qwen2.5-Coder series instruct models on the BigCodeBench-Instruct. As shown in Table 13, the Qwen2.5-Coder-7B-Instruct outperformed other instruct models with similar parameter sizes, achieving higher accuracy scores on both the full and hard subsets, reaching 41.0% on the full subset and 18.2% on the hard subset, demonstrating the Qwen2.5-Coder series instruct models’ powerful code generation capabilities.
@@
To better demonstrate our model’s effectiveness on real-world competitive programming tasks, we conduct evaluation of the Qwen-2.5-Coder series instruct models on the LiveCodeBench (2305-2409) dataset. As illustrated in Table 13, the Qwen-2.5-Coder-7B-Instruct model achieved an impressive Pass@1 accuracy of 37.6%, significantly outperforming other models of comparable parameter scales. Notably, it also surpassed larger models such as CodeStral-22B and DS-Coder-33B-Instruct, underscoring the Qwen-2.5-Coder series’ exceptional capabilities in handling complex code generation challenges.
@@
As demonstrated by the evaluation results in Table 14, Qwen2.5-Coder-7B-Instruct consistently outperforms other models with the same number of parameters, including DS-Coder-V2-Lite-Instruct, in code generation tasks across eight programming languages. With an average accuracy of 76.5%, Qwen2.5-Coder-7B-Instruct surpasses even larger models such as CodeStral-22B and DS-Coder-33B-Instruct (despite having over 20 billion parameters), highlighting its powerful code generation capabilities in multiple programming languages.
Figure 6: The McEval Performance of Qwen2.5-Coder-7B-Instruct compared with popular open-source large code models with similar size.
@@
To comprehensively assess the code generation capabilities of the Qwen2.5-Coder series models across a broader range of programming languages, we evaluated them on the McEval benchmark (Chai et al., 2024), which spans 40 programming languages and includes 16,000 test cases. As shown in Figure 6, the Qwen2.5-Coder-7B-Instruct model excels when compared to other open-source models on the McEval benchmark, particularly across a wide range of programming languages. Its average accuracy not only exceeds that of much larger models such as DS-Coder-33B-Instruct and CodeStral-22B but also demonstrates a notable advantage over models of comparable parameter size.
@@
DS-Coder-1.3B-Instruct 1.3B 14.8 28.1

### v2

@@ 6.2 Code Completion
CodeStral-22B 22B 76.7 82.5 86.0 82.7
DS-Coder-33B-Base 33B 80.1 89.0 86.8 86.2
Qwen2.5-Coder-32B 32B 81.5 91.0 89.4 88.3
Table 7: Performance of different approaches on the Humaneval-FIM Tasks. Average refers to a weighted mean calculated based on the number of samples for each language.
In real-world scenarios, code completion often depends on accessing cross-file context and dependencies. CrossCodeEval is a benchmark that requires a deep understanding of this cross-file context to accurately complete the code. In our evaluation, we set a maximum sequence length of 8192 tokens, designate a maximum output length of 50 tokens, and impose a limit of 2048 tokens for the cross-file context. For the cross-file context, we use the official BM25 search results provided by Ding et al. (2024). We evaluate performance using Exact Match (EM) and Edit Similarity (ES) metrics. Table 8 shows that the Qwen2.5-Coder-32B achieves state-of-the-art performance with a 3.7% improvement. Qwen2.5-Coder outperforms all the models with a comparable model size. Meanwhile, Qwen2.5-Coder-7B has a comparable performance with other models exceeding 20 billion parameters.
Model Python Java TypeScript C# Average
EM ES EM ES EM ES EM ES EM ES
0.5B+ Models
Qwen2.5-Coder-0.5B 22.7 66.2 21.7 66.8 21.9 67.2 32.1 75.4 24.6 68.9
1B+ Models
DS-Coder-1.3B-Base 33.4 72.6 34.9 74.5 36.7 76.4 46.6 83.5 37.9 76.8
Qwen2.5-Coder-1.5B 35.5 74.3 37.9 76.5 37.6 77.4 49.8 84.5 40.2 78.2
3B+ Models
StarCoder2-3B 11.0 62.7 11.6 69.7 8.8 75.8 8.2 71.2 9.9 69.8
Qwen2.5-Coder-3B 38.4 76.1 42.8 79.8 41.6 80.5 56.7 87.1 44.9 80.9
6B+ Models
StarCoder2-7B 10.9 63.1 8.3 71.0 6.7 76.8 7.3 72.1 8.3 70.8
DS-Coder-6.7B-Base 41.1 79.2 39.9 80.1 46.3 82.4 55.0 86.9 45.6 82.1
DS-Coder-V2-Lite-Base 41.8 78.3 46.1 81.2 44.6 81.4 58.7 87.9 47.8 82.2
CodeQwen1.5-7B 40.7 77.8 47.0 81.6 45.8 82.2 59.7 87.6 48.3 82.3
Qwen2.5-Coder-7B 42.4 78.6 48.1 82.6 46.8 83.4 59.7 87.9 49.3 83.1
14B+ Models
StarCoder2-15B 28.2 70.5 26.7 71.0 24.7 76.3 25.2 74.2 26.2 73.0
Qwen2.5-Coder-14B 47.7 81.7 54.7 85.7 52.9 86.0 66.4 91.1 55.4 86.1
20B+ Models
CodeStral-22B 49.3 82.7 44.1 71.1 51.0 85.0 53.7 83.6 49.5 80.6
DS-Coder-33B-Base 44.2 80.4 46.5 82.7 49.2 84.0 55.2 87.8 48.8 83.7
Qwen2.5-Coder-32B 49.2 82.1 56.4 86.6 54.9 87.0 68.0 91.6 57.1 86.8
Table 8: Performance of different approaches on the CrossCodeEval Tasks.
CrossCodeLongEval is a long context benchmark on cross file code completion tasks. In our evaluation, we set a maximum sequence length of 8192 tokens and set the maximum output as 256 tokens for function completion and 50 tokens for other tasks. The cross-file context is truncated to 2048 tokens. For the cross-file context, we use the official BM25 search results provided by Wu et al. (2024a). We evaluate performance using Exact Match (EM) and Edit Similarity (ES) metrics. Qwen2.5-Coder-32B achieves state-of-the-art performance, as detailed in Table 9. The Qwen2.5-Coder series surpasses all other models of a similar size. All models demonstrate low Exact Match (EM) results on function completion tasks, likely due to the complexity of generating multi-line code snippets that are challenging to match precisely.
Model Chunk Completion Function completion Average
EM ES EM ES EM ES
0.5B+ Models
Qwen2.5-Coder-0.5B 29.8 64.2 9.5 38.0 19.7 51.1
1B+ Models
DS-Coder-1.3B-Base 40.6 71.9 9.6 39.4 25.1 55.7
Qwen2.5-Coder-1.5B 44.2 73.9 12.4 44.4 28.3 59.2
3B+ Models
StarCoder2-3B 18.5 62.0 10.2 39.2 14.3 50.6
Qwen2.5-Coder-3B 46.6 76.1 13.5 46.4 30.0 61.3
6B+ Models
StarCoder2-7B 19.4 63.6 10.2 40.0 14.8 51.8
DS-Coder-6.7B-Base 48.4 78.2 10.7 42.4 29.6 60.3
DS-Coder-V2-Lite-Base 49.5 77.1 11.4 43.1 30.4 60.1
CodeQwen1.5-7B 48.2 77.5 6.4 30.6 27.3 54.1
Qwen2.5-Coder-7B 52.4 79.3 14.4 48.4 33.4 63.8
14B+ Models
StarCoder2-15B 21.3 53.7 7.8 30.5 14.6 42.1
Qwen2.5-Coder-14B 56.9 81.8 15.4 49.8 36.1 65.8
20B+ Models
CodeStral-22B 56.7 81.8 10.5 37.8 33.6 59.8
DS-Coder-33B-Base 52.0 79.9 11.9 44.3 32.0 62.1
Qwen2.5-Coder-32B 57.3 82.1 16.4 50.8 36.9 66.4
Table 9: Performance of different approaches on the CrossCodeLongEval Tasks.
RepoEval is a benchmark designed to evaluate repository-level code completion capabilities across three granularities: line, API invocation, and function body completion. In our evaluation, we set a maximum sequence length of 8192 tokens, set the maximum output as 256 tokens for function completion and 50 tokens for other tasks, and impose a limit of 2048 tokens for the cross-file context. Besides, we utilize the official sparse retriever (Lu et al., 2022) to extract the cross-file context. We evaluate performance using Exact Match (EM) and Edit Similarity (ES) metrics. As shown in Table 10, Qwen2.5-Coder-32B achieves state-of-the-art performance with an average improvement of 7.9% EM and 4.2% ES compared to DS-Coder-33B-Base. Furthermore, Qwen2.5-Coder-14B and Qwen2.5-Coder-7B achieve comparable performance to models with more than 20B parameters, while maintaining state-of-the-art results among models of similar size.
Model Line Function API Average
EM ES EM ES EM ES EM ES
0.5B+ Models
Qwen2.5-Coder-0.5B 44.2 72.6 4.6 48.0 35.6 68.5 28.1 63.0
1B+ Models
DS-Coder-1.3B-Base 58.7 80.4 6.1 48.8 45.8 75.0 36.9 68.1
Qwen2.5-Coder-1.5B 59.8 82.6 10.6 52.4 51.0 80.1 40.5 71.7
3B+ Models
StarCoder2-3B 22.3 67.4 3.1 51.6 20.6 70.1 15.3 63.0
Qwen2.5-Coder-3B 64.9 85.0 12.3 55.8 54.7 81.3 44.0 74.0
6B+ Models
StarCoder2-7B 19.5 67.6 4.0 53.5 19.1 72.8 14.2 64.7
DS-Coder-6.7B-Base 63.1 85.5 9.9 53.3 52.3 81.7 41.7 73.5
DS-Coder-V2-Lite-Base 66.5 85.4 10.8 53.9 53.1 81.3 43.4 73.5
CodeQwen1.5-7B 59.7 81.5 4.8 44.3 46.1 77.5 36.9 67.8
Qwen2.5-Coder-7B 67.3 86.1 13.2 55.2 58.4 83.9 46.3 75.1
14B+ Models
StarCoder2-15B 30.9 62.5 5.5 43.7 21.7 60.3 19.4 55.5
Qwen2.5-Coder-14B 74.3 90.1 14.1 59.5 63.4 87.3 50.6 79.0
20B+ Models
Codestral-22B-v0.1 40.9 51.7 9.9 49.2 24.8 40.8 25.2 47.2
DS-Coder-33B-Base 66.5 86.6 10.3 52.9 54.2 83.5 43.7 74.3
Qwen2.5-Coder-32B 76.1 90.5 13.6 57.5 65.1 87.6 51.6 78.5
Table 10: Performance of different approaches on the RepoEval Tasks.
SAFIM is a syntax-aware fill-in-the-middle benchmark that emphasizes AST-based code completion, specifically targeting algorithmic blocks, control-flow expressions, and API function calls. The benchmark consists of 17,720 examples from 8,590 code files created after April 2022, deliberately avoiding overlap with mainstream pretraining corpora. For evaluation, we use pass@1 rate as the metric for algorithmic and control-flow tasks, and Exact Match (EM) for API completion tasks.
@@ 6.3 Code Reasoning
CodeQwen1.5-7B 7B 44.8 40.1
Qwen2.5-Coder-7B 7B 56.5 56.0
14B+ Models
StarCoder2-15B 15B 46.1 47.6
Qwen2.5-Coder-14B 14B 60.6 66.4
@@ 6.4 Math Reasoning
Mathematics and coding have always been closely intertwined. Mathematics forms the foundational discipline for coding, while coding serves as a vital tool in mathematical fields. As such, we expect an open and powerful code model to exhibit strong mathematical capabilities as well. To assess Qwen2.5-Coder’s mathematical performance, we selected five popular benchmarks, including MATH (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), MMLU-STEM (Hendrycks et al., 2020) and TheoremQA (Chen et al., 2023). Table 12 highlights Qwen2.5-Coder’s strengths in mathematics, which likely stem from two key factors: first, the model’s strong foundation built on Qwen2.5, and second, the careful mixing of code and mathematical data during training, which has ensured a well-balanced performance across these domains.
@@
CodeQwen1.5-7B 7B 10.6 37.7 39.6 15.8
Qwen2.5-Coder-7B 7B 46.6 83.9 67.6 34.0
14B+ Models
StarCoder2-15B 15B 23.7 57.7 49.2 20.5
Qwen2.5-Coder-14B 14B 52.8 88.7 73.9 39.6
@@
DS-Coder-V2-Base 21/236B 50.6 85.8 76.0 39.4
Qwen2.5-Coder-32B 32B 57.2 91.1 75.1 43.1
Table 12: Performance of various models on four math benchmarks, named MATH, GSM8K, MMLU STEM and TheoremQA respectively.
@@
Base Pro Redux
0.5B+ Models
Qwen2.5-Coder-0.5B 0.5B 42.0 13.3 40.6
@@
DS-Coder-1.3B-Base 1.3B 25.8 11.4 24.5
Qwen2.5-Coder-1.5B 1.5B 53.6 23.1 50.9
3B+ Models
StarCoder2-3B 3B 36.6 15.5 37.0
Qwen2.5-Coder-3B 3B 61.2 32.0 59.5
@@
StarCoder2-7B 7B 38.8 17.2 38.6
DS-Coder-6.7B-Base 6.7B 36.4 16.7 36.5
DS-Coder-V2-Lite-Base 2.4/16B 60.5 33.4 58.3
CodeQwen1.5-7B 7B 40.5 17.2 41.2
Qwen2.5-Coder-7B 7B 68.0 40.1 66.6
14B+ Models
StarCoder2-15B 15B 64.1 24.3 48.8
Qwen2.5-Coder-14B 14B 75.2 49.3 72.4
@@
DS-Coder-33B-Base 33B 39.4 18.4 38.7
Qwen2.5-Coder-32B 32B 79.1 50.4 77.5
@@
DS-Coder-6.7B-Base 6.7B 36.4 40.2 57.6 53.8
DS-Coder-V2-Lite-Base 2.4/16B 57.3 38.8 72.9 76.1
CodeQwen1.5-7B 7B 35.7 42.2 59.8 56.0
Qwen2.5-Coder-7B 7B 60.9 50.6 72.9 76.8
14B+ Models
@@
Qwen2.5-Coder-14B 14B 66.0 55.2 76.8 80.2
@@
DS-Coder-V2-Base 21/236B 64.3 41.4 83.7 86.0
Qwen2.5-Coder-32B 32B 70.5 54.2 80.8 83.0
Table 14: General performance of different models on four popular general benchmarks, ARC-Challenge, TruthfulQA, WinoGrande and HellaSwag.
@@ 6.6 Long-Context Evaluation
Long context capability is crucial for code LLMs, serving as the core skill for understanding repository-level code and becoming a code agent. However, most of the current code models still have very limited support for length, which hinders their potential for practical application. Qwen2.5-Coder aims to further advance the progress of open-source code models in long context modeling. To achieve this, we have collected and constructed long sequence code data at the repository level for pre-training. Through careful data proportioning and organization, we have enabled it to support input lengths of up to 128K tokens.
@@
Figure 6: The long context ability of Qwen2.5-Coder, evaluated by Needle in the Code.
@@ 7 Evaluation on Instruct Models
Qwen2.5-Coder-0.5B-Instruct https://hf.co/Qwen/Qwen2.5-Coder-0.5B-Instruct
Qwen2.5-Coder-1.5B-Instruct https://hf.co/Qwen/Qwen2.5-Coder-1.5B-Instruct
Qwen2.5-Coder-3B-Instruct https://hf.co/Qwen/Qwen2.5-Coder-3B-Instruct
Qwen2.5-Coder-7B-Instruct https://hf.co/Qwen/Qwen2.5-Coder-7B-Instruct
Qwen2.5-Coder-14B-Instruct https://hf.co/Qwen/Qwen2.5-Coder-14B-Instruct
Qwen2.5-Coder-32B-Instruct https://hf.co/Qwen/Qwen2.5-Coder-32B-Instruct
CodeQwen1.5-7B-Chat https://hf.co/Qwen/CodeQwen1.5-7B-Chat
CodeLlama-7B-Instruct https://hf.co/meta-llama/CodeLlama-7b-Instruct-hf
CodeLlama-13B-Instruct https://hf.co/meta-llama/CodeLlama-13b-Instruct-hf
CodeLlama-34B-Instruct https://hf.co/meta-llama/CodeLlama-34b-Instruct-hf
CodeLlama-70B-Instruct https://hf.co/meta-llama/CodeLlama-70b-Instruct-hf
@@
Starcoder2-15B-Instruct-v0.1 https://hf.co/bigcode/starcoder2-15b-instruct-v0.1
CodeStral-22B-v0.1 https://hf.co/mistralai/Codestral-22B-v0.1
Yi-Coder-1.5B-Chat https://hf.co/01-ai/Yi-Coder-1.5B-Chat
Yi-Coder-9B-Chat https://hf.co/01-ai/Yi-Coder-9B-Chat
@@ 7.1 Code Generation
We also assessed the code generation capabilities of the Qwen2.5-Coder series instruction models using the EvalPlus (Liu et al., 2023) dataset. As shown by the results in Table 16, our Qwen2.5-Coder-7B-Instruct model demonstrated exceptional accuracy, significantly outperforming other models with a comparable parameter count. Remarkably, it even surpassed larger models with over 20 billion parameters, such as CodeStral-22B and DS-Coder-33B-Instruct. Furthermore, our Qwen2.5-Coder-32B-Instruct model achieved the highest performance on EvalPlus, even outperforming DS-Coder-V2-Instruct, making it the most powerful open-source code model to date.
@@
DS-Coder-1.3B-Instruct 1.3B 65.9 60.4 65.3 54.8 22.8 3.4 5.1
Yi-Coder-1.5B-Chat 1.5B 69.5 64.0 65.9 57.7 23.8 11.5 4.8
Qwen2.5-Coder-1.5B-Instruct 1.5B 70.7 66.5 69.2 59.4 32.5 6.8 6.1
3B+ Models
Qwen2.5-Coder-3B-Instruct 3B 84.1 80.5 73.6 62.4 35.8 14.2 10.8
@@
CodeLlama-7B-Instruct 7B 40.9 33.5 54.0 44.4 21.9 3.4 7.1
DS-Coder-6.7B-Instruct 6.7B 74.4 71.3 74.9 65.6 35.5 10.1 15.5
CodeQwen1.5-7B-Chat 7B 83.5 78.7 77.7 67.2 39.6 18.9 7.9
Yi-Coder-9B-Chat 9B 82.3 74.4 82.0 69.0 38.1 11.5 17.2
DS-Coder-V2-Lite-Instruct 2.4/16B 81.1 75.6 82.8 70.4 36.8 16.2 16.3
Qwen2.5-Coder-7B-Instruct 7B 88.4 84.1 83.5 71.7 40.4 18.2 18.2
13B+ Models
CodeLlama-13B-Instruct 13B 40.2 32.3 60.3 51.1 28.5 9.5 6.1
Starcoder2-15B-Instruct-v0.1 15B 67.7 60.4 78.0 65.1 37.2 11.5 12.1
Qwen2.5-Coder-14B-Instruct 14B 70.7 66.5 69.2 59.4 32.5 6.8 23.4
@@
CodeLlama-34B-Instruct 34B 48.2 40.2 61.1 50.5 29.0 8.8 8.4
CodeStral-22B-v0.1 22B 81.1 73.2 78.2 62.2 41.8 16.9 23.7
DS-Coder-33B-Instruct 33B 81.1 75.0 80.4 70.1 42.0 17.6 21.3
CodeLlama-70B-Instruct 70B 72.0 65.9 77.8 64.6 40.7 11.5 3.3
DS-Coder-V2-Instruct 21/236B 85.4 82.3 89.4 75.1 48.2 24.3 27.9
Qwen2.5-Coder-32B-Instruct 32B 92.7 87.2 90.2 75.1 49.6 27.0 31.4
Closed-APIs
Claude-3.5-Sonnet-20240620 - 89.0 81.1 87.6 72.0 45.3 25.7 32.1
Claude-3.5-Sonnet-$2$ 92.1 86.0 91.0 74.6 45.3 23.6 31.6
GPT-4o-mini-2024-07-18 - 87.8 84.8 86.0 72.2 46.9 23.6 28.3
GPT-4o-2024-08-06 - 92.1 86.0 86.8 72.5 50.1 25.0 34.6
o1-mini - 97.6 90.2 93.9 78.3 46.3 23.0 60.0
o1-preview - 95.1 88.4 93.4 77.8 49.3 27.7 43.1
@@
13B+ Models
CodeLlama-13B-Instruct 13B 42.7 40.5 42.2 24.0 39.0 - 32.3 13.9 -
Starcoder2-15B-Instruct-v0.1 15B 68.9 53.8 50.9 62.7 57.9 59.6 53.4 24.7 54.0
Qwen2.5-Coder-14B-Instruct 14B 89.0 79.7 85.1 84.2 86.8 84.5 80.1 47.5 79.6
@@
CodeLlama-70B-Instruct 70B 67.8 58.2 53.4 36.7 39.0 - 58.4 29.7 -
DS-Coder-V2-Instruct 21/236B 90.2 82.3 84.8 82.3 83.0 84.5 79.5 52.5 79.9
Qwen2.5-Coder-32B-Instruct 32B 92.7 80.4 79.5 82.9 86.8 85.7 78.9 48.1 79.4
Closed-APIs
Claude-3.5-Sonnet-20240620 - 89.6 86.1 82.6 85.4 84.3 84.5 80.7 48.1 80.2
Claude-3.5-Sonnet-$2$ 93.9 86.7 88.2 87.3 88.1 91.3 82.6 52.5 83.8
GPT-4o-mini-2024-07-18 - 87.2 75.9 77.6 79.7 79.2 81.4 75.2 43.7 75.0
GPT-4o-2024-08-06 - 90.9 83.5 76.4 81.0 83.6 90.1 78.9 48.1 79.1
o1-mini - 95.7 90.5 93.8 77.2 91.2 92.5 84.5 55.1 85.1
o1-preview - 96.3 88.0 91.9 84.2 90.6 93.8 90.1 47.5 85.3
Table 17: The performance of different models on instruct format MultiPL-E.
@@
The instruct split provided by BigCodeBench (Zhuo et al., 2024) is designed to evaluate the code generation capabilities of instruction-based models. We evaluated the Qwen2.5-Coder series instruct models on the BigCodeBench-Instruct dataset. As indicated in Table 16, the Qwen2.5-Coder-7B-Instruct model outperformed other instruct models with comparable parameter sizes, achieving notably high accuracy scores on both the full and hard subsets, reaching 41.0% on the full subset and 18.2% on the hard subset. This highlights the robust code generation capabilities of the Qwen2.5-Coder instruct models. Furthermore, the Qwen2.5-Coder-32B-Instruct achieved accuracy rates of 49.6% on the complete split and 27.0% on the hard split, establishing it as the best-performing open-source code generation model and surpassing several closed-source APIs.
@@
To further demonstrate our model’s effectiveness on real-world competitive programming tasks, we evaluated the Qwen-2.5-Coder series instruct models on the LiveCodeBench (2407-2409) dataset. As shown in Table 16, the Qwen-2.5-Coder-7B-Instruct model achieved an impressive Pass@1 accuracy of 37.6%, significantly outperforming other models with similar parameter counts. Notably, it also outperformed larger models, such as CodeStral-22B-v0.1 and DS-Coder-33B-Instruct. Additionally, our Qwen-2.5-Coder-32B-Instruct model achieved an accuracy of 31.4%, surpassing all open-source code generation models and reaching a level comparable to many closed-source APIs.
@@
As shown by the evaluation results in Table 17, Qwen2.5-Coder-7B-Instruct consistently outperforms other models with similar parameter counts, such as DS-Coder-V2-Lite-Instruct, in code generation tasks across eight programming languages. Both Qwen2.5-Coder-7B-Instruct and Qwen2.5-Coder-14B-Instruct even surpass larger models, like CodeStral-22B and DS-Coder-33B-Instruct (which have over 20 billion parameters), underscoring their strong code generation capabilities across multiple languages. Our Qwen2.5-Coder-32B-Instruct model achieves comparable performance to the DS-Coder-V2-Instruct model with only 32 billion parameters, bringing it very close to the performance of several closed-source APIs.
Figure 7: The McEval Performance of Qwen2.5-Coder-32B-Instruct compared with popular open-source large code models with similar size.
Figure 8: The MdEval Performance of Qwen2.5-Coder-32B-Instruct compared with popular open-source large code models with similar size.
@@
To comprehensively assess the code generation capabilities of the Qwen2.5-Coder series models across a broader range of programming languages, we evaluated them on the McEval benchmark (Chai et al., 2024), which spans 40 programming languages and includes 16,000 test cases. As shown in Figure 7, the Qwen2.5-Coder-32B-Instruct model excels when compared to other open-source models on the McEval benchmark, particularly across a wide range of programming languages.
@@
DS-Coder-1.3B-Instruct 1.3B 12.9 28.1
Yi-Coder-1.5B-Chat 1.5B 19.9 24.9
@@
13B+ Models
CodeLlama-13B-Instruct 13B 47.5 41.1
Starcoder2-15B-Instruct-v0.1 15B 45.5 50.9
Qwen2.5-Coder-14B-Instruct 14B 69.5 79.5
@@
Qwen2.5-Coder-32B-Instruct 32B 75.2 83.4
Closed-APIs
Claude-3.5-Sonnet-20240620 - 75.5 81.8
Claude-3.5-Sonnet-$2$ 84.4 87.2
GPT-4o-mini-2024-07-18 - 67.5 78.4
GPT-4o-2024-08-06 - 78.6 89.2
o1-mini - 91.6 96.2
o1-preview - 86.5 81.4
@@
MdEval
Qwen2.5-Coder is further evaluated on the comprehensive multilingual code debugging benchmark MdEval (Liu et al., 2024b) across 18 languages. Compared to the multilingual code generation benchmark McEval (Chai et al., 2024), MdEval provides the buggy code with example test cases (1.2K samples) to LLM for generating the correct code. Figure 8 demonstrates that the Qwen2.5-Coder-32B-Instruct achieves a comparable or better performance even compared to LLMs with larger model sizes.
Human Preference Alignment
To evaluate the alignment performance of Qwen2.5-Coder-32B-Instruct with the human preferences, we adopted an internal annotated evaluation benchmark called CodeArena, including nearly 400 human-curated samples. Similar to Chatbot Arena (Chiang et al., 2024), we use CodeArena to emulate user code-related prompts in realistic environments. We use GPT-4o as the evaluation model for preference alignment, employing an “A vs. B win” evaluation method, which measures the percentage of instances in the test set where the score of A exceeds the score of B. The results in Figure 9 demonstrate the advantage of Qwen2.5-Coder-32B-Instruct in preference alignment.
Figure 9: The CodeArena Performance of Qwen2.5-Coder-32B-Instruct compared with popular open-source large code models with similar size.